the length of a rectangle is 3 times its width. the perimeter of the rectangle is 72 cm. find the demensions of the retangle.

Answer by checkley79(3341)

P=2L+2W
L=3W
72=2*3W+2W
72=6W+2W
72=8W
W=72/8
W=9 ANS. FOR THE WIDTH.
L=3*9=27 ANS. FOR THE LENGTH.
PROOF.
72=2*27+2*9
72=54+18
72=72